Problem 10-14 Basic Variance Analysis [LO10-1, LO10-2, LO10-3]

Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.20 ounces $ 25.00 per ounce $ 55.00
Direct labor 0.50 hours $ 15.00 per hour 7.50
Variable manufacturing overhead 0.50 hours $ 3.00 per hour 1.50
Total standard cost per unit $ 64.00

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 12,000 ounces at a cost of $282,000.
  2. There was no beginning inventory of materials; however, at the end of the month, 2,750 ounces of material remained in ending inventory.

  3. The company employs 25 lab technicians to work on the production of Fludex. During November, they each worked an average of 110 hours at an average pay rate of $11.50 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $2,400.

  5. During November, the company produced 4,100 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 25 technicians employed in the production of Fludex consisted of 5 senior technicians and 20 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1

Solution 1a:

Standard quantity of material for actual production = 4100*2.20 = 9020 ounce

Actual quantity of material purchased = 12000 ounce

Actual quantity of material used = 12000 - 2750 = 9250 ounce

Standard price of material = $25 per ounce

Actual price of material = $282,000 / 12000 = $23.50

Material price variance = (SP - AP) * AQ purchased = ($25 - $23.50) * 12000 = $18,000 F

Material quantity variance = (SQ - AQ) * SR = (9020 - 9250) * $25 = $5,750 U

Solution 1b:

As price offered by the new supplier is lesser than standard price of material, therefore company should sign long term purchase contract with the new supplier.

Solution 2a:

Standard hours of direct labor = 4100*0.5 = 2050 hours

Standard rate of direct labor = $15 per hour

Actual hours of direct labor = 25*110 = 2750 hours

Actual rate of direct labor = $11.50 per hour

Direct labor rate variance = (SR - AR) * AH = ($15 - $11.50) * 2750 = $9,625 F

Direct labor efficiency variance = (SH - AH) * SR = (2050 - 2750) * $15 = $10,500 U

Solution 2b:

Employing more assistant rather senior technician resulted in favorable direct labor rate variance but unfavorable laor efficiency variance. Further unfavorable efficiency variance is higher than favorable rate variance, therefore it is recommended new labor mix should not be continued.

Solution 3:

Standard hours of direct labor = 4100*0.5 = 2050 hours

Standard rate of variable overhead= $3 per hour

Actual hours of direct labor = 2750

Actual rate of variable overhead = $2,400 / 2750 = $0.872727

Variable overhead rate variance = (SR - AR) * AH = ($3 - $0.872727) * 2750 = 5,850 F

Variable overhead efficiency variance = (SH - AH) * SR = (2050 - 2750) * $3 = $2,100 U